OpenStudy (aroub):

If you want to solve system of linear equation by substitution you ( for example ) : Equation (1) Equation (2) You find x in equation (1) and then substitute it in equation (2) and when you find y in equation (2) you substitute it in equation (1) ??

OpenStudy (angela210793):

Aroub :O:O

OpenStudy (anonymous):

:)(:

OpenStudy (diyadiya):

Yes! :D =D

OpenStudy (aroub):

lol, sounds confusing right? -.- Can you solve this one for me? ( with steps ) PLEASE =) 8x+4y=12 10x+5y=15

OpenStudy (accessdenied):

8x + 4y = 12 10x + 5y = 15 2x + y = 3 2x = -y + 3 x = -1/2 y + 3/2 10(-1/2 y + 3/2) + 5y = 15 -5y + 15 + 5y = 15 6 = 15 << no solution

OpenStudy (diyadiya):

both equations are same

OpenStudy (accessdenied):

oops, i half-updated some mistake should be 15=15

OpenStudy (aroub):

Oh I got that too! I thought I did something wrong :P lol thank youuu all :DD

OpenStudy (diyadiya):

8x+4y=12 is same as 2x+y=12 10x+5y=15 is same as 2x+y=12 So infinitely many solutions

OpenStudy (aroub):

Wait! It's no solution or infinite solution? O.o

OpenStudy (accessdenied):

infinitely many -- it'd be no-solutions if the numbers weren't equal (in the case I had before I changed to 15=15. x.x)

OpenStudy (aroub):

Aooh! Thank you :D